# MIMO Detection with Spatial Sigma-Delta ADCs: A Variational Bayesian Approach

### v1

@@ V Extension to MIMO Detection with $2^{\rm nd}$-Order $\Sigma\Delta$ Quantization
V-A Proposed SD-VB Approach for Data Detection
@@
The mean-field VB inference method aims to optimize the variational distribution $q(\mathbf{s},\mathbf{r},\gamma)$ such that
@@
$\hskip 20.00003pt+\ln p(y_{i}|r_{i})+\ln p(r_{i+1}|r_{i},y_{i},r_{i-1},y_{i-1},\mathbf{s},\gamma;\mathbf{H})$
$\hskip 20.00003pt+\ln p(r_{i+2}|r_{i+1},y_{i+1},r_{i},y_{i},\mathbf{s},\gamma;\mathbf{H})\big{\rangle}\big{\}},$ (35)
which is expanded as in (V-A) on the next page,
@@
$\hskip 10.00002pt-2\mathrm{e}^{-j\phi}(r_{i}-y_{i})+\mathrm{e}^{-j2\phi}(r_{i-1}-y_{i-1})|^{2}\big{\rangle}-\big{\langle}\gamma|r_{i+2}-\mathbf{h}_{i+2}^{T}\mathbf{s}-2\mathrm{e}^{-j\phi}(r_{i+1}-y_{i+1})+\mathrm{e}^{-j2\phi}(r_{i}-y_{i})|^{2}\big{\rangle}\big{\}}$
@@
$\hskip 10.00002pt+\mathrm{e}^{j\phi}\mathbf{h}_{i+1}^{T}\mathbf{s}-\mathrm{e}^{-j\phi}(r_{i-1}-y_{i-1})|^{2}\big{\rangle}-\langle\gamma\rangle\big{\langle}|r_{i}-y_{i}+\mathrm{e}^{j2\phi}r_{i+2}-\mathrm{e}^{j2\phi}\mathbf{h}_{i+2}^{T}\mathbf{s}-2\mathrm{e}^{j\phi}(r_{i+1}-y_{i+1})|^{2}\big{\rangle}\big{\}}$
@@
×$\;\times\mathcal{CN}\big{(}r_{N-1};(c_{N-1}+2d_{N-1})/5,1/(5\langle\gamma\rangle)\big{)},$ (37)
@@
$v_{i}$ =$=\left\{\begin{array}[]{ll}\frac{c_{i}+2d_{i}+f_{i}}{6}=\langle r_{i}\rangle-\frac{u_{i}-2\mathrm{e}^{j\phi}u_{i+1}+\mathrm{e}^{j2\phi}u_{i+2}}{6}&\textrm{ for }i=1,\ldots,N-2,\\ \frac{c_{N-1}+2d_{N-1}}{5}=\langle r_{N-1}\rangle-\frac{u_{N-1}-2\mathrm{e}^{j\phi}u_{N}}{5}&\textrm{ for }i=N-1,\\ c_{N}=\langle r_{N}\rangle-u_{N}&\textrm{ for }i=N.\\ \end{array}\right.$ (43)
@@
$\hskip 20.00003pt+\sum_{i=1}^{N}\ln p(r_{i}|r_{i-1},y_{i-1},r_{i-2},y_{i-2},\mathbf{s},\gamma;\mathbf{H})\bigg{\rangle}\bigg{\}},$
@@
$\hskip 20.00003pt-2\mathrm{e}^{-j\phi}(r_{i-1}-y_{i-1})+\mathrm{e}^{-j2\phi}(r_{i-2}-y_{i-2})|^{2}\big{\rangle}\bigg{\}}$
@@
which is expanded as in (V-A) on the next page,
@@
The proposed VB method for MIMO detection with $2^{\rm nd}$-order $\Sigma\Delta$ quantization is summarized in Algorithm 2, where the parameter $\epsilon$ is the numerator of $\hat{\gamma}$ in the variational distributions of (V-A), (V-A), and (38) for updating $r_{i}$, $r_{N-1}$, and $r_{N}$, respectively.
@@
6:            $+\mathrm{Tr}\{\mathbf{H}\bm{\Sigma}_{\mathbf{s}}\mathbf{H}^{H}\}\big{)}$
@@
19:        $u_{i+2}\leftarrow u_{i+2}-\mathrm{e}^{-j2\phi}(\hat{r}_{i}^{t}-\hat{r}_{i}^{t+1})$ only for $i<N-1$
@@
V-B Computational Complexity Analysis
@@
TABLE I: Algorithm Computational Complexity
@@
The computational complexity of the proposed algorithms is analyzed here assuming $N\geq K$. For iterative algorithms, the complexity order is given per iteration. For LMMSE, the computation of $\mathbf{H}\bm{\Sigma}_{\mathbf{s}}\mathbf{H}^{H}$ for the linear detector in (13) requires a complexity of $\mathcal{O}(NK^{2})$, while calculating $\mathbf{U}^{-1}\bm{\Sigma}_{\mathbf{q}}\mathbf{U}^{-H}$ requires $\mathcal{O}(N^{3})$ operations due to the matrix inversion. In addition, the denoiser for $K$ users contributes a complexity of $\mathcal{O}(|\mathcal{S}|K)$. Thus, the overall computational burden of the LMMSE detector is dominated by the matrix inversion, resulting in an overall complexity of order $\mathcal{O}(N^{3}+|\mathcal{S}|K)$. For the $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms in Algs. $1$ and $2$, respectively, the cost to compute $\mathrm{Tr}\{\mathbf{H}\bm{\Sigma}_{\mathbf{s}}\mathbf{H}^{H}\}=\sum_{k=1}^{K}\tau_{s_{k}}\|\bar{\mathbf{h}}_{k}\|^{2}$ in step 5 is $\mathcal{O}(NK)$, which dominates that for computing the residual vector $\|\mathbf{u}\|^{2}$ which is $\mathcal{O}(N)$. The computation in step 18 of Algorithm 1 and step 22 of Algorithm 2 adds a complexity of $\mathcal{O}(|\mathcal{S}|)$ for each user. The overall per-iteration complexity of Algs. $1$ and $2$ is thus $\mathcal{O}(NK+|\mathcal{S}|K)$, and we see that the $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms have the same complexity as that for the MF-VB and MF-QVB approaches, although the $\Sigma\Delta$ methods can provide a significant improvement in MIMO detection performance under the given assumptions of spatial oversampling or sectored users. Note that the complexities of the VB-based algorithms presented in Table I also indicate the number of iterations $T$ required for convergence. Our simulations indicate that $T$ is typically less than $10$.
VI Numerical Results
@@
In this section, we present illustrative numerical results for the performance of the proposed $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms compared with state-of-the-art data detection methods such as the matched-filter quantized VB (MF-QVB) in [10] and LMMSE-based detection for various scenarios. We implement all VB-based algorithms with a maximum of $50$ iterations and consider scenarios with $100$ transmitted data symbols. The noise variance $N_{0}$ is set based on the SNR, which is defined as
${\rm SNR}=\frac{\mathbb{E}[\|\mathbf{Hs}\|^{2}]}{\mathbb{E}[\|\mathbf{n}\|^{2}]}=\frac{K}{NN_{0}}.$ (49)
@@
Unless otherwise stated, all cases assume the number of paths as $L=20$, the width of the angular sector as $\Theta=40^{\circ}$ and assume it is centered at $\theta_{0}=0^{\circ}$. We assume all users lie within the same azimuth angular range, with AoAs drawn uniformly from the interval $[-20^{\circ},20^{\circ}]$. We also assume the number of BS antennas is $N=128$, the number of users is $K=16$, the phase shift of the $\Sigma\Delta$ array is $\phi=2\pi\frac{d}{\lambda}\sin(\theta_{0})$, $d=\lambda/6$, and the users employ QPSK signaling.
@@
    The LMMSE receiver presented in Section III implemented with the $1^{\rm st}$-order $\Sigma\Delta$ architecture and $1$-bit quantizers,
@@
Figure 5: SER performance of various algorithms vs. SNR with 1-bit and 3-bit quantizers, $\theta_{0}=20^{\circ}$, $\Theta=40^{\circ}$, and $d=\lambda/6$.
@@
Fig. 5 shows the data detection performance improvement of the algorithms using 1-bit and 3-bit quantization. The SD-VB algorithms outperform MF-QVB since the $\Sigma\Delta$ receivers shape the quantization noise away from the spatial frequencies of interest. The SD-VB algorithms also significantly outperform the LMMSE-based receiver because the latter’s reliance on linear processing makes it ineffective at separating users with highly correlated channels. We note that the SERs of all algorithms that process quantized signals converge to their error floors at high SNR because the quantization noise becomes the dominant source of error.
@@
The performance of $2^{\rm nd}$-order SD-VB significantly improves when the ADC resolution is increased from $1$ to $3$ bits, achieving the best SER among the considered algorithms, while MF-QVB performs the worst. This advantage is due to the more aggressive noise shaping of the $2^{\rm nd}$-order $\Sigma\Delta$ architecture, greatly reducing the quantization noise in lower spatial frequency ranges. Although MF-QVB shows some improvement with $3$-bit quantization, it remains inefficient for cases with users confined to an angular sector since the quantization noise is uniformly spread across all spatial frequencies. While the MF-VB algorithm employs ideal/infinite quantization, it also experiences an error floor due to the channel correlation[28]. The $2^{\rm nd}$-order SD-VB with only $3$-bit quantizers outperforms MF-VB since it implicitly exploits knowledge of the angular sector to which the users are confined and the resulting spatial correlation. While LMMSE receiver provided in (13) is designed for $1$-bit quantization, it shows considerable SER improvement when the quantization resolution is increased to $3$ bits. Thus, in subsequent simulations we will only show results for the $3$-bit LMMSE implementation.
@@
Figure 6: SER performance vs. azimuth angular spread $\Theta$, with 3-bit quantizers, SNR $=5$ dB, $\theta_{0}=0^{\circ}$, and ${d=\lambda/6}$.
@@
Figure 7: SER performance vs. $d/\lambda$, with 3-bit quantizers, $N=128$, $\Theta=40^{\circ}$, $\theta_{0}=0^{\circ}$, and ${\rm SNR}=5$ dB.
@@
Fig. 7 presents the effect of antenna spacing and wavelength on the SER performance of all detection algorithms for an array with a fixed number of antennas ($N=128$). Observations similar to those in the previous example can be made here. In particular, when $d$ is very small, the array aperture is significantly reduced and none of the methods are able to counteract the extreme channel correlation that results from the narrow angular sector of $\Theta=40^{\circ}$. The SER of all algorithms improves with increasing $d$, although the $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms provide the best performance for values of $d$ around $1/4$ to $1/3$, and degrade for larger $d$ since the benefit of oversampling is lost. As the antenna spacing increases, the reduced channel correlation benefits MF-VB and MF-QVB.
@@
Figure 8: SER performance vs. ($d/\lambda$, $N$), with a fixed array $d_{0}=14\,$cm, 3-bit quantizers, $K=16$, $\Theta=40^{\circ}$, $\theta_{0}=0^{\circ}$, and ${\rm SNR}=8$ dB.
In Fig. 8 we also study the impact of the antenna spacing $d$, but in this case we fix the array aperture $d_{0}=Nd$ so that as $d$ decreases, the number of antennas $N$ increases. For example, if the aperture is set at $d_{0}=Nd=14$ cm, the wavelength is $\lambda=c/f=6.7$ mm at $f=45$ GHz, and as $d$ decreases from $\lambda/2$ to $\lambda/14$, $N$ increases from $42$ to $294$. As shown in the figure, in this case there is much less degradation for small $d$ since the larger array aperture maintains a more constant level of channel correlation. Interestingly, the LMMSE receiver benefits the most from the increase in $N$, and yields better performance than the MF-VB and SD-VB for very small $d$. In this case, the best trade-off between the number of observations and the channel correlation for the SD-VB algorithms occurs for $d=\lambda/6$. It is critical to note that the two previous examples have ignored the impact of mutual coupling, which will become important as $d$ becomes much smaller than $\lambda/2$. Interestingly, as shown in [41], the $\Sigma\Delta$ approach benefits to some extent from mutual coupling due to the increase in correlation between the signals at adjacent antennas. This is an issue that requires further study for the VB-based algorithms.
@@
Figure 9: SER performance vs. number of users with 3-bit quantizers, ${\rm SNR}=8$ dB, $N=80$, $d=\lambda/6$, $\theta_{0}=0^{\circ}$, and $\Theta=40^{\circ}$.
@@
In Fig. 9, we show the effects of the number of users $K$ on the detection performance. When $K$ decreases, the lower spatial channel correlation enhances the data detection performance of all algorithms. The $2^{\rm nd}$-order SD-VB algorithm achieves the lowest SER among the considered algorithms when $K\leq 10$, including MF-VB with infinite precision ADCs. On the other hand, MF-VB has the best performance for large $K$, but in these cases the SER is already quite high. MF-QVB performs the worst due to its use of matched filtering to handle multiuser interference, which is suboptimal with high channel correlation, even with relatively few users. LMMSE excels with a small number of users since it is able to more effectively eliminate inter-user interference, achieving a lower SER than MF-VB, $1^{\rm st}$-order SD-VB, and MF-QVB when $K=10$.
@@
Figure 10: SER performance vs. number of antennas with 3-bit quantizers, SNR$=8$ dB, $K=30$, $d=\lambda/6$, $\theta_{0}=0^{\circ}$, and $\Theta=40^{\circ}$.
@@
In Fig. 10 we present results for data detection with different numbers of antennas. As $N$ increases, the $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms show significant SER improvement, while the improvement is marginal for LMMSE and MF-QVB. Moreover, the SER performance gap between $2^{\rm nd}$-order SD-VB and the other detection algorithms becomes wider as the number of antennas increases.
@@
Figure 11: SER performance vs. $\Sigma\Delta$ ADC resolution $b$ bits with SNR$=12$ dB, $\theta_{0}=0$, $d=\lambda/6$, and $\Theta=40^{\circ}$.
@@
Figure 12: SER performance vs. different steering angles $\phi=2\pi\frac{d}{\lambda}\sin(\theta)$ of $\Sigma\Delta$ ADCs with 3-bit quantizers, SNR$=12$ dB and $d=\lambda/6$.

### v2

@@ IV Extension to MIMO Detection with $2^{\rm nd}$-Order $\Sigma\Delta$ Quantization
IV-A Proposed SD-VB Approach for Data Detection
@@
This conditional probability follows a complex Gaussian distribution, driven by the input-output relationship of the $2^{\rm nd}$-order $\Sigma\Delta$ approach in (38), where the fact that the noise $n_{i}$ is Gaussian implies that (41) is Gaussian. The mean-field VB inference method aims to optimize the variational distribution $q(\mathbf{s},\mathbf{r},\gamma)$ such that
@@
$\qquad+\ln p(y_{i}|r_{i})+\ln p(r_{i+1}|r_{i},y_{i},r_{i-1},y_{i-1},\mathbf{s},\gamma;\mathbf{H})$
$\qquad+\ln p(r_{i+2}|r_{i+1},y_{i+1},r_{i},y_{i},\mathbf{s},\gamma;\mathbf{H})\big\rangle\big\},$ (43)
which is expanded as in (IV-A) on the next page,
@@
$\quad-2\mathrm{e}^{-j\phi}(r_{i}-y_{i})+\mathrm{e}^{-j2\phi}(r_{i-1}-y_{i-1})|^{2}\big\rangle-\big\langle\gamma|r_{i+2}-\mathbf{h}_{i+2}^{T}\mathbf{s}-2\mathrm{e}^{-j\phi}(r_{i+1}-y_{i+1})+\mathrm{e}^{-j2\phi}(r_{i}-y_{i})|^{2}\big\rangle\big\}$
@@
$\quad+\mathrm{e}^{j\phi}\mathbf{h}_{i+1}^{T}\mathbf{s}-\mathrm{e}^{-j\phi}(r_{i-1}-y_{i-1})|^{2}\big\rangle-\langle\gamma\rangle\big\langle|r_{i}-y_{i}+\mathrm{e}^{j2\phi}r_{i+2}-\mathrm{e}^{j2\phi}\mathbf{h}_{i+2}^{T}\mathbf{s}-2\mathrm{e}^{j\phi}(r_{i+1}-y_{i+1})|^{2}\big\rangle\big\}$
@@
$\;\times\mathcal{CN}\big(r_{N-1};(c_{N-1}+2d_{N-1})/5,1/(5\langle\gamma\rangle)\big),$ (45)
@@
[…]
@@
$\qquad+\sum_{i=1}^{N}\ln p(r_{i}|r_{i-1},y_{i-1},r_{i-2},y_{i-2},\mathbf{s},\gamma;\mathbf{H})\bigg\rangle\bigg\},$
@@
$\qquad-2\mathrm{e}^{-j\phi}(r_{i-1}-y_{i-1})+\mathrm{e}^{-j2\phi}(r_{i-2}-y_{i-2})|^{2}\big\rangle\bigg\}$
@@
which is expanded as in (IV-A) on the next page,
@@
The proposed VB method for MIMO detection with $2^{\rm nd}$-order $\Sigma\Delta$ quantization is summarized in Algorithm 2, where the parameter $\epsilon$ is the numerator of $\hat{\gamma}$ in the variational distributions of (IV-A), (IV-A), and (46) for updating $r_{i}$, $r_{N-1}$, and $r_{N}$, respectively. We refer this algorithm as the $2^{\rm nd}$-order SD-VB algorithm.
@@
6:                        $\mathrm{Tr}\{\mathbf{H}\bm{\Sigma}_{\mathbf{s}}\mathbf{H}^{H}\}\big)$
@@
19:    $u_{i+2}\leftarrow u_{i+2}-\mathrm{e}^{-j2\phi}(\hat{r}_{i}^{t}-\hat{r}_{i}^{t+1})$ only for i<N1i<N-1
@@
Remark 3: The $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms are designed to address the nonlinear quantization effects of spatial $\Sigma\Delta$ ADCs in massive MIMO systems, particularly for discrete constellations such as PSK and QAM in angularly sectored mmWave channels. The SD-VB algorithms are not optimal for Gaussian input signals because they approximate the posterior of the discrete latent variables and rely on mean-field factorization, which does not fully capture the continuous statistical properties of Gaussian signals.
The $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms, based on CAVI, factorize $r_{i}$, $s_{k}$, and $\gamma$, iteratively updating $q(s_{k})$, $q(r_{i})$, and $q(\gamma)$ until they converge to a locally optimal solution. This efficient approach suits massive MIMO systems with highly non-linear SD quantization, where local optima yield acceptable SER. We will show through simulations in Section V that these algorithms achieve low SER, outperforming state-of-the-art methods, especially with optimized antenna spacing, with the $2^{\rm nd}$-order SD-VB enhancing detection by reaching a better local optimum under various scenarios and mutual coupling effects.
IV-B Computational Complexity Analysis
@@
TABLE II: Algorithm Computational Complexity
@@
The computational complexity of the proposed algorithms is analyzed here assuming $N\geq K$. For iterative algorithms, the complexity order is given per iteration. For LMMSE, the computation of $\mathbf{H}\bm{\Sigma}_{\mathbf{s}}\mathbf{H}^{H}$ for the linear detector in (16) requires a complexity of $\mathcal{O}(NK^{2})$, while calculating $\mathbf{U}^{-1}\bm{\Sigma}_{\mathbf{q}}\mathbf{U}^{-H}$ requires $\mathcal{O}(N^{3})$ operations due to the matrix inversion. In addition, the denoiser for $K$ users contributes a complexity of $\mathcal{O}(|\mathcal{S}|K)$. Thus, the overall computational burden of the LMMSE detector is dominated by the matrix inversion, resulting in an overall complexity of order $\mathcal{O}(N^{3}+|\mathcal{S}|K)$. For the $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms in Algs. $1$ and $2$, respectively, the cost to compute $\mathrm{Tr}\{\mathbf{H}\bm{\Sigma}_{\mathbf{s}}\mathbf{H}^{H}\}=\sum_{k=1}^{K}\tau_{s_{k}}\|\bar{\mathbf{h}}_{k}\|^{2}$ in step 5 is $\mathcal{O}(NK)$, which dominates that for computing the residual vector $\|\mathbf{u}\|^{2}$ which is $\mathcal{O}(N)$. The computation in step 18 of Algorithm 1 and step 22 of Algorithm 2 adds a complexity of $\mathcal{O}(|\mathcal{S}|)$ for each user. The overall per-iteration complexity of Algs. $1$ and $2$ is thus $\mathcal{O}(NK+|\mathcal{S}|K)$, and we see that the $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms have the same complexity as that for the MF-VB and MF-QVB approaches, although the $\Sigma\Delta$ methods can provide a significant improvement in MIMO detection performance under the given assumptions of spatial oversampling or sectored users. Note that the complexity of the VB-based algorithms listed in Table I also reflects the number of iterations $T$ needed for convergence. Numerical simulations suggest that $T$ is typically below $20$.
V Numerical Results
@@
In this section, we present illustrative numerical results for the performance of the proposed $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms compared with state-of-the-art data detection methods such as the matched-filter quantized VB (MF-QVB) in [10] and LMMSE-based detection for various scenarios. We implement all VB-based algorithms with a maximum of $50$ iterations and consider scenarios with $100$ transmitted data symbols. The channel matrix $\mathbf{H}$ is normalized such that each column has unit norm, i.e., $\mathbb{E}\big[\|\bar{\mathbf{h}}_{k}\|^{2}\big]=1,\forall k$ [33]. The SNR is defined as
${\rm SNR}=\frac{\mathbb{E}\big[\|\mathbf{Hs}\|^{2}\big]}{\mathbb{E}\big[\|\mathbf{n}_{c}\|^{2}\big]}=\frac{K}{\mathrm{Tr}\{\mathbf{R}\}}.$ (57)
@@
If the system is without mutual coupling, the noise covariance matrix $\mathbf{R}$ reduces to $N_{0}\mathbf{I}$, and the SNR becomes ${\rm SNR}=K/(NN_{0})$}. Unless otherwise stated, all cases assume the number of paths as $L=20$, the width of the angular sector as $\Theta=40^{\circ}$ and assume it is centered at $\theta_{0}=0^{\circ}$. We assume all users lie within the same azimuth angular range, with AoAs drawn uniformly from the interval $20^{\circ},20^{\circ}]$. We also assume the number of BS antennas is $N=128$, the number of users is $K=16$, the phase shift of the $\Sigma\Delta$ array is $\phi=2\pi\frac{d}{\lambda}\sin(\theta_{0})$, $d=\lambda/6$, and the users employ QPSK signaling. Taking the mutual coupling effect into account, we set the circuit parameters in (6) and (II-A1) as $\sigma_{i}^{2}=2kTB/R$, and $\sigma_{u}^{2}=2kTBR$, leading to the noise ratio $R_{N}=\sqrt{\sigma_{u}^{2}/\sigma_{i}^{2}}=R$ where $R=50~\Omega$, $T=290~\mathrm{K}$, $\rho=0$, and $B=20~\mathrm{MHz}$. The factor of 2 in the relationship $\sigma_{n}^{2}=2kTBR$ appears because we are accounting for noise in both the antennas and the LNAs.
@@
    The LMMSE receiver presented in Section II-C implemented with the $1^{\rm st}$-order $\Sigma\Delta$ architecture and 11-bit quantizers,
@@
We first study the convergence of all iterative algorithms with $3$-bit quantization, as shown in Fig. 6. All algorithms are observed to converge in less than 20 iterations, which demonstrates the effectiveness of using CAVI in our VB framework.
Figure 6: Convergence of the iterative algorithms with SNR=12=12 dB, $\theta_{0}=20^{\circ}$, $d=\lambda/6$, $\Theta=40^{\circ}$, 33-bit quantization, and QPSK signaling. The plots are obtained by averaging over 200 trials.
Figure 7: SER performance vs. SNR with $\theta_{0}=20^{\circ}$, $\Theta=40^{\circ}$, $d=\lambda/4$, QPSK, and without mutual coupling. For the quantized system, solid lines represent results with 3-bit ADCs; dashed lines represent results with 1-bit ADCs.
Figure 8: SER performance vs. SNR with 3-bit quantizer, $\theta_{0}=20^{\circ}$, $\Theta=40^{\circ}$, $d=\lambda/4$, $16$-QAM. Solid lines represent results without mutual coupling; dashed lines represent results with mutual coupling effects.
@@
We show the data detection performance improvement of the proposed SD-VB algorithms using 1-bit and 3-bit ADCs with QPSK in Fig. 7, and using 3-bit ADCs with $16$-QAM and MC effects in Fig. 8. The SD-VB algorithms outperform MF-QVB for both 1-bit and 3-bit ADCs since the $\Sigma\Delta$ receivers shape the quantization noise away from the spatial frequencies of interest. The SD-VB algorithms also significantly outperform the LMMSE-based receiver because the latter’s reliance on linear processing makes it ineffective at separating users with highly correlated channels. The $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms with $1$-bit quantization and QPSK exhibit higher SER than MF-VB since the latter operates as an unquantized (full-precision) system. We note that coarse quantization (e.g., 1-bit or 3-bit ADCs) introduces quantization noise that cannot be reduced and persists even at high SNR. As a result, the symbol error rates (SERs) of all algorithms processing quantized signals converge to their error floors at high SNR, since quantization noise becomes the dominant source of error.
@@
The performance of $2^{\rm nd}$-order SD-VB significantly improves when the ADC resolution is increased from $1$ to $3$ bits, achieving the best SER among the considered algorithms, while MF-QVB performs the worst. This advantage is due to the more aggressive noise shaping of the $2^{\rm nd}$-order $\Sigma\Delta$ architecture, greatly reducing the quantization noise in lower spatial frequency ranges. Although MF-QVB shows some improvement with $3$-bit quantization for both QPSK and 16-QAM, it remains inefficient for cases with users confined to an angular sector since the quantization noise is uniformly spread across all spatial frequencies. While the MF-VB algorithm employs ideal/infinite quantization, it also experiences an error floor due to the channel correlation [33]. The $2^{\rm nd}$-order SD-VB with only $3$-bit quantizers outperforms MF-VB since it implicitly exploits knowledge of the angular sector to which the users are confined and the resulting spatial correlation. While the LMMSE receiver provided in (16) is designed for $1$-bit quantization, it shows considerable SER improvement when the quantization resolution is increased to $3$ bits. Thus, in subsequent simulations we will only show results for the $3$-bit LMMSE implementation.
The effect of MC is shown in Fig. 8 for 16-QAM signaling, illustrating that the SER performance of LMMSE and MF-QVB degrades considerably with MC, while the proposed SD-VB algorithms have only a slight decrease at high SNRs. This is due to two reasons. First, since spatial $\Sigma\Delta$ techniques rely on the correlation between signals at adjacent antennas, the correlation introduced by MC actually provides some benefit for $\Sigma\Delta$ that counteracts the corresponding loss due to channel correlation. In addition, when MC introduces colored noise, the $\Sigma\Delta$ noise-shaping effect treats it similarly to quantization noise, pushing it outside the signal bandwidth. The LMMSE detector is implemented using (17), which incorporates the known colored noise covariance matrix $\mathbf{R}$ induced by MC. Despite this, the LMMSE approach still exhibits SER degradation compared to SD-VB. This is due to the limitations of the Bussgang decomposition, which assumes Gaussian inputs, and the approximation of the quantization noise covariance $\bm{\Sigma}_{\mathbf{q}}^{\rm{MC}}$ as diagonal [52, 50].
@@
Figure 9: SER performance vs. azimuth angular spread $\Theta$, with 3-bit quantizers, SNR =5=5 dB, $\theta_{0}=0^{\circ}$, $d=\lambda/6$}, and QPSK.
@@
Figure 10: SER performance vs. $d/\lambda$, with 3-bit quantizers, $N=128$, $\Theta=40^{\circ}$, $\theta_{0}=0^{\circ}$, QPSK, ${\rm SNR}=5$ dB, and without MC.
@@
Fig. 10 presents the effect of antenna spacing and wavelength on the SER performance of all detection algorithms for an array with a fixed number of antennas ($N=128$) without considering MC. Observations similar to those in the previous example can be made here. In particular, when $d$ is very small, the array aperture is significantly reduced and none of the methods are able to counteract the extreme channel correlation that results from the narrow angular sector of $\Theta=40^{\circ}$. The SER of all algorithms improves with increasing $d$, although the $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms provide the best performance for values of $d$ around $1/4$ to $1/3$, and degrade for larger $d$ since the benefit of oversampling is lost. As the antenna spacing increases, the reduced channel correlation benefits MF-VB and MF-QVB. Although reducing the antenna spacing to around $1/3$ to $1/4$ wavelength can enhance the SER for the SD-VB algorithms, it reduces the achievable data rate due to the reduced array aperture [52, 18]. Similarly, lower quantization resolutions (e.g., 1-bit or 3-bit) improve power efficiency but limit the amount of information per sample. As studied in [18], spectral efficiency in $\Sigma\Delta$ MIMO systems scales with the oversampling rate, highlighting a trade-off between noise reduction and capacity. In future works, we will analyze the capacity of MIMO systems with $2^{\rm nd}$-order $\Sigma\Delta$ ADCs, leveraging the probabilistic modeling capabilities of the VB framework.
@@
Figure 11: SER performance vs. ($d/\lambda$, NN), with mutual coupling effects, a fixed array $d_{0}=14\,$cm, 3-bit quantizers, QPSK, $K=16$, $\Theta=40^{\circ}$, $\theta_{0}=0^{\circ}$, and ${\rm SNR}=18$ dB.
In Fig. 11 we study the impact of the antenna spacing $d$ in the presence of MC, but in this case, we fix the array aperture $d_{0}=Nd$ so that as $d$ decreases, the number of antennas $N$ increases. For example, if the aperture is set at $d_{0}=Nd=14$ cm, the wavelength is $\lambda=c/f=6.7$ mm at $f=45$ GHz, and as $d$ decreases from $\lambda/2$ to $\lambda/14$, $N$ increases from $42$ to $294$. As shown in the figure, in this case there is much less degradation for small $d$ since the larger array aperture maintains a more constant level of channel correlation. The SD-VB algorithms maintain superior SER performance over the other algorithms under MC due to their ability to exploit the MC-induced spatial correlation, and because their noise-shaping process treats colored noise introduced by MC similarly to quantization noise, pushing it away from the spatial frequencies of interest. In this case, the best trade-off between the number of observations and the channel correlation for the SD-VB algorithms occurs for $d=\lambda/6$. LMMSE struggles with high channel correlation under MC since its linear processing cannot effectively de-correlate the users, leading to higher SER. The MF-VB benefits from infinite precision but is still limited by channel correlation, performing worse than the SD-VB algorithms when $d>\lambda/6$ and for narrow sectors. The spatial $\Sigma\Delta$ architecture achieves its best performance with small antenna spacings and confined angular sectors but this potentially limits its real-world application. Nonetheless, these constraints enable compact, power-efficient array designs, e.g., an aperture of 14 cm can fit $126$ elements at $d=\lambda/6$.
@@
Figure 12: SER performance vs. number of users with 3-bit quantizers, ${\rm SNR}=8$ dB, QPSK, $N=80$, $d=\lambda/6$, $\theta_{0}=0^{\circ}$, and $\Theta=40^{\circ}$.
@@
In Fig. 12, we show the effects of the number of users $K$ on the SER performance. When $K$ decreases, the lower spatial channel correlation enhances the data detection performance of all algorithms. The $2^{\rm nd}$-order SD-VB algorithm achieves the lowest SER among the considered algorithms when $K\leq 20$, including MF-VB with infinite precision ADCs. On the other hand, MF-VB has the best performance for large $K$, but in these cases the SER is already quite high. MF-QVB performs the worst due to its use of matched filtering to handle multiuser interference, which is suboptimal with high channel correlation, even with relatively few users. LMMSE excels with a small number of users since it is able to more effectively eliminate inter-user interference, achieving a lower SER than MF-VB, $1^{\rm st}$-order SD-VB, and MF-QVB when $K=10$.
@@
Figure 13: SER performance vs. number of antennas with 3-bit quantizers, SNR=8=8 dB, QPSK, $K=30$, $d=\lambda/6$, $\theta_{0}=0^{\circ}$, and $\Theta=40^{\circ}$.
@@
In Fig. 13 we present results for data detection with different numbers of antennas. As $N$ increases, the $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms show significant SER improvement, while the improvement is marginal for LMMSE and MF-QVB. The SER performance gap between $2^{\rm nd}$-order SD-VB and the other detection algorithms becomes wider as the number of antennas increases. Note that MF-VB outperforms $1^{\rm st}$-order SD-VB since it is implemented under the assumption of zero quantization noise.
@@
Figure 14: SER performance vs. $\Sigma\Delta$ ADC resolution $b$ bits with SNR=12=12 dB, QPSK, $\theta_{0}=0$, $d=\lambda/6$, and $\Theta=40^{\circ}$.
@@
Figure 15: SER performance vs. steering angles $\phi=2\pi\frac{d}{\lambda}\sin(\theta)$ of $\Sigma\Delta$ ADCs with 3-bit quantizers, SNR at $12$ dB, QPSK, and $d=\lambda/6$.